# Boosting Graph Pooling with Persistent Homology

### v1

@@ 1 Introduction
Persistent homology (PH) is a powerful tool in the field of topological data analysis, which is capable of evaluating stable topological invariant properties from unstructured data in a multi-resolution fashion (Edelsbrunner & Harer, 2022). Concretely, PH derives an increasing sequence of simplicial complex subsets by applying a filtration function (see Fig. 1(a)). According to the fact that PH is at least as expressive as Weisfeiler-Lehman (WL) hierarchy (Horn et al., 2021), there recently emerged a series of works seeking to merge PH into graph neural networks (GNNs), delivering competitive performance on specific tasks (Zhao et al., 2020; Horn et al., 2021). Standard schemes of existing works achieve this by employing pre-calculated topological features (Zhao et al., 2020) or placing learnable filtration functions in the neural architectures (Hofer et al., 2020; Horn et al., 2021). Such integration of PH features is claimed to enable GNNs to emphasize persistent topological sub-structures. However, it is still unclear to what extent the feature-level integration of PH is appropriate and how to empower GNNs with PH other than utilizing features.
@@
Graph pooling (GP) in parallel plays an important role in a series of graph learning methods (Grattarola et al., 2022), which hierarchically aggregates an upper-level graph into a more compact lower-level graph. Typically, GP relies on calculating an assignment matrix taking into account local structural properties such as community (Müller, 2023) and cuts (Bianchi et al., 2020). Though the pooling paradigm in convolutional neural networks (CNNs) is quite successful (Krizhevsky et al., 2012), some researchers raise concerns about its effectiveness and applicability in graphs. For example, Mesquita et al. (2020) challenges the local-preserving usage of GP by demonstrating that random pooling even leads to similar performance. Till now, it remains opaque what property should be preserved for pooled topology to better facilitate the downstream tasks.
@@
From Fig. 1(a), it is readily observed that PH and GP both seek to coarsen/sparsify a given graph in a hierarchical fashion: while PH gradually derives persistent sub-topology (substructures that have meaningful topology) by adjusting the filtering parameter, GP obtains a sub-graph by performing a more aggressive cut-off. In a sense of understanding a graph through a hierarchical lens, PH and GP turn out to align with each other well.
@@
Driven by this observation, in this paper, we investigate the mechanism of aligning PH and GP so as to mutually reinforce each other. To this end, we conduct experiments by running a pioneer GP method DiffPool (Ying et al., 2018) to conduct graph classification on several datasets and at the same time use the technique in (Hofer et al., 2020) to compute PH information. We manually change pooling ratio and see what proportion of meaningful topological information (characterized by the ratio of non-zero persistence) is naturally preserved at the final training stage. Surprisingly, the correspondence is quite stable regardless of different datasets (see Fig. 1(b)), which implies the monotone trend between the pooling ratio and non-zero persistence is commonly shared by a large range of graph data. As a consequence, we develop a natural way to integrate PH and GP in both feature and topology levels. Concretely, in addition to concatenating vectorized PH diagram as supplementary features, we further enforce the coarsened graph to preserve topological information as much as possible with a specially designed PH-inspired loss function. Hence we term our method Topology-Invariant Pooling (TIP). TIP can be flexibly injected into a variety of existing GP methods, and demonstrates a consistent ability to provide substantial improvement over them. We summarize our contributions as follows:
@@ 2 Related work
Graph pooling. Graph pooling has been used in various applications, which can reduce the graph size while preserving its structural information. Early methods are based on clustering to coarsen graphs, such as the greedy clustering method Graclus (Dhillon et al., 2007), non-negative matrix factorization of the adjacency matrix (Bacciu & Di Sotto, 2019), and spectral clustering (Ma et al., 2019). Recently, learnable graph pooling methods have gained popularity, which learn to select important nodes in an end-to-end manner. DiffPool (Ying et al., 2018) follows a hierarchical learning structure by utilizing GNNs to learn clusters and gradually aggregate nodes into a coarser graph. MinCutPool (Bianchi et al., 2020) optimizes a normalized cut objective to partition graphs into clusters. DMoNPool (Müller, 2023) optimizes the modularity of graphs to ensure high-quality clusters. SEP (Wu et al., 2022) generates clusters in different hierarchies simultaneously without compromising local structures. These methods are classified as dense pooling due to the space complexity they incur. Despite their effectiveness, dense pooling methods have been criticized for high memory cost and complexity (Cangea et al., 2018). Therefore, various sparse pooling methods have been proposed, such as Top-K (Gao & Ji, 2019), ASAPool (Ranjan et al., 2020), and SAGPool (Lee et al., 2019). These methods coarsen graphs by selecting a subset of nodes based on a ranking score. As they drop some nodes in the pooling process, these methods are criticized for their limited capacity to retain essential information, with potential effects on the expressiveness of preceding GNN layers (Bianchi & Lachi, 2023).
Persistent homology in GNNs. PH is a technique to calculate topological features of structured data, and many approaches have been proposed to use PH in graph machine learning due to the high expressiveness of topological features on graphs (Hofer et al., 2017). Since non-isomorphic graphs may exhibit different topological features, the combination of PH and the Weisfeiler-Lehman (WL) algorithm leads to stronger expressive power (Rieck et al., 2019). This encourages further exploration on equipping GNNs with topological features. Zhao et al. (2020) propose that message passing in GNNs can be effectively reweighted using topological features. Hofer et al. (2020) and Horn et al. (2021) provide theoretical and practical insights that filtraions in PH can be purely learnable, enabling flexible usage of topological features in GNNs. However, existing methods tend to view PH merely as a tool for providing supplementary information to GNNs, resulting in only marginal improvements and limited interpretability.
@@ 3 Background
Let $\mathcal{G}=(V,E)$ be an undirected graph with $n$ nodes and $m$ edges, where $V$ and $E$ are the node and the edge sets, respectively. Nodes in attributed graphs are associated with features, and we denote by $V=\{(v,\mathbf{x}_{v})\}_{v\in 1:n}$ the set of nodes $v$ with $d$ dimensional attribute $\mathbf{x}_{v}$. It is also practical to represent the graph with an adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ and the node feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$
@@
We focus on the general message-passing GNN framework that updates node representations by iteratively aggregating information from neighbors (Gilmer et al., 2017). Concretely, the $k$-th layer of such GNNs can be expressed as:
@@
where $\mathbf{\theta}^{(k)}$ is the trainable parameter, and $\mathrm{M}$ is the message propagation function. Numbers of $\mathrm{M}$ have been proposed in previous research (Kipf & Welling, 2016; Hamilton et al., 2017). A complete GNN is typically instantiated by stacking multiple layers of Eq. 1. Hereafter we denote by $\mathrm{GNN}(\cdot)$ an arbitrary such multi-layer GNN for brevity.
@@
GP in GNNs is a special layer designated to produce a coarsened or sparsified sub-graph. Formally, GP can be formulated as $\mathcal{G}\mapsto\mathcal{G}_{P}=(V_{P},E_{P})$ such that the number of nodes $|V_{P}|\leq n$. GP layers can be placed into GNNs in a hierarchical fashion to persistently coarsen the graph. Typical GP approaches (Ying et al., 2018; Bianchi et al., 2020; Müller, 2023) rely on learning a soft cluster assignment matrix $\mathbf{S}^{(l)}\in\mathbb{R}^{n_{l-1}\times n_{l}}$:
@@
These approaches differ from each other in the way to produce $\mathbf{S}$, which is used to inject a bias in the formation of clusters. In our work, we select three GP methods, i.e., DiffPool (Ying et al., 2018), MinCutPool (Bianchi et al., 2020), and DMoNPool (Müller, 2023), to cope with. Details of these methods are in Appendix A.
@@
Simplicial complexes stand as the focal point within the realm of algebraic topology. An assembly of simplices of certain dimensions constitutes a simplicial complex denoted as $K$. A graph can be seen as a low-dimensional simplicial complex that only contains 0-simplices (vertices) and 1-simplices (edges) (Horn et al., 2021). The simplest kind of topological features describing graphs are Betti numbers, formally denoted as $\beta_{0}$ for the number of connected components and $\beta_{1}$ for the number of cycles.
Despite the limited expressive power of these two numbers, it can be improved by evaluating them alongside a filtration. Filtrations are scalar-valued functions of the form $f:V\cup E\rightarrow\mathbb{R}$ , which assigns each vertice and edge a value. Changes in the Betti numbers, named as persistent Betti numbers, can subsequently be monitored throughout the progress of the filtration: by considering a threshold $(a\in\mathbb{R})$, we can analyze the subgraph originating from the pre-image of $((-\infty,a])$ of $f$, denoted as $(f^{-1}((-\infty,a]))$. The image of $f$ leads to a set of node values $a_{1}<\cdots<a_{n}$ and generates a sequence of nested subgraphs of the form $\emptyset\subseteq\mathcal{G}_{0}\subseteq\ldots\mathcal{G}_{k}\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$, where $\mathcal{G}_{k}=(V_{k},E_{k})$ is a subgraph of $\mathcal{G}$ with $V_{k}:=\left\{v\in V\mid f\left(\mathbf{x}_{v}\right)\leq a_{k}\right\}$ and $E_{k}:=\left\{(v,w)\in E\mid\max\left\{f\left(x_{v}\right),f\left(x_{w}\right)\right\}\leq a_{k}\right\}$. This process is also known as persistent homology (denoted as $ph(\cdot)$) on graphs. Typically, persistent Betti numbers are characterized in a persistence diagram (PD) as $\mathrm{ph}(\mathcal{G},f)=\{\mathcal{D}_{0},\mathcal{D}_{1},...\}$, which is made up of tuples $(a_{i},a_{j})\in\mathbb{R}^{2}$, with $a_{i}$ and $a_{j}$ representing the creation and destruction of a topological feature respectively (see Fig. 1(c)). The absolute difference in function values $|a_{j}-a_{i}|$ is called the persistence of a topological feature, where high persistence corresponds to features of the function, while low persistence is typically considered as noise (Horn et al., 2021; Rieck, 2023). Moreover, we follow the settings in previous works (Horn et al., 2021; Hofer et al., 2017) to extend $\mathcal{D}_{1}$ as follows: (1) each cycle is paired with the edge that created it; (2) edges $e$ that do not create a cycle (still in this circle) are assigned a ‘dummy’ tuple value, such as $(f(e),f(e))$; (3) all other edges will be paired with the maximum value of the filtration. Therefore, $\mathcal{D}_{1}$ consists of as many tuples as the number of edges $m$.
@@ 4.1 Overview
Figure 2: Overview of our method, where the shaded part corresponds to one layer of Topology-Invariant Pooling.
@@ 4.2 Topology-Invariant Pooling
In many real-world applications, such as molecular graphs analysis (Swenson et al., 2020; Hofer et al., 2020), topological features of graphs are of utmost importance. However, typical GNNs fail to capture certain topological structures in graphs, such as cycles (Bouritsas et al., 2022; You et al., 2021; Huang et al., 2022). Moreover, in dense graph pooling, graphs are pooled without preserving any topology. Even if we manage to make GNN topology-aware, the coarsened graph has no meaningful topology at all, impairing the use of GNNs in these tasks. To overcome these limitations, we propose to inject topological information into GP. We resort to PH to characterize the importance of edges. Note that for those edges do not form cycles, their creation and destruction are the same, leading to zero persistence.
The core of PH is the notion of filtration, which presents a challenging task to choose the right filtration. As the coarsened graph evolves in each training step, integrating PH into GP demands multiple computations of filtrations. To address this, we place learnable filtration (LF) functions for incorporating PH information, which is flexible and efficient as demonstrated by Hofer et al. (2020). LF relies on node features and graph topology, which are readily available in GP. Consequently, LF can be seamlessly integrated into GP with minimal computational overhead. Specifically, we employ an MLP network $\Phi(\cdot)$ as the filtration function together with $\mathrm{sigmoid}(\cdot)$ to map node features $\mathbf{X}\in\mathbb{R}^{n\times d}$ into $n$ scalar values. Recently, an increasing amount of attention has been devoted to cycles (Bouritsas et al., 2022; You et al., 2021; Huang et al., 2022) due to their significant relevance to downstream tasks in various domains such as biology (Koyutürk et al., 2004), chemistry (Murray & Rees, 2009), and social network analysis (Jiang et al., 2010). Recognizing that cycles offer an intuitive representation of graph structure, we focus on the one-dimensional PD associated with cycles. Following the standard way in GP (Eq. 2 3 4), we additionally propose the subsequent modules to inject PH into GP at both feature and topology levels.
@@
One major limitation of utilizing LF is that the computation process is unconscious of edge weights, i.e. edges with non-negative weights will be treated equally, so PH cannot directly extract meaningful topological features from $\mathbf{A}^{(l)}$. Besides, rethinking GP in Eq. 3, the coarsened adjacency matrix has limited expressive power for two reasons. First, although $\mathbf{S}^{(l)}$ is a soft assignment matrix obtained by $\mathrm{softmax}(\cdot)$, each element still has nonzero values, i.e. $\mathbf{A}^{(l)}$ is always densely connected. Second, the edge weights may span a wide range by multiplication (see Appendix D for empirical evidence). These drawbacks hinder the stability and generalization power of the subsequent message passing layers (Gong & Cheng, 2019). None of the existing GP methods can handle these problems properly.
@@
where $\mathbf{A}^{(l)}$ is first normalized in the range of $[0,1]$, and $\mathrm{resample}(\cdot)$ is performed independently for each matrix entry using the Gumbel-softmax trick (Jang et al., 2016). In practice, only the upper triangular matrix is resampled to make it symmetric and we add self-loops to the graph.
@@
The aforementioned mechanism can explicitly inject topological information into graphs, but it relies on the condition that the coarsened graph retains certain essential sub-topology. To this end, we propose an additional loss function to guide the GP process. Intuitively, the coarsened graph should exhibit similarity to the original graph in terms of topology. Since the computation of PH is differentiable, one possible approach is to directly minimize the differences between the PDs of the original graph and the coarsened graph. However, this implementation would require computing the Wasserstein distance between two PDs through optimal transport (Yan et al., 2022), which is intractable in training due to its complexity. Considering that our objective is to estimate the difference, we instead propose vectorizing the PDs and minimizing their high-order statistical features (Okabe et al., 2018). Specifically, we use several transformations (denoted as $\mathrm{transform(\cdot)}$) and concatenate the output, including the triangle point transformation, the Gaussian point transformation and the line point transformation introduced in (Carrière et al., 2020) to convert the tuples in PD into vector $\mathbf{h}_{t}$ $(t\in[1,m])$. We calculate the mean vector $\mathbf{\mu}$ as well as the second-order statistics as the standard deviation vector $\mathbf{\sigma}$ as:
$\mathbf{h}_{t}=\mathrm{transform(\mathcal{D}_{1})}\qquad\qquad\qquad$ (7)

### v2

@@ 1 Introduction
Persistent homology (PH) is a powerful tool in the field of topological data analysis, which is capable of evaluating stable topological invariant properties from unstructured data in a multi-resolution fashion [8]. Concretely, PH derives an increasing sequence of simplicial complex subsets by applying a filtration function (see Fig. 1(a)). According to the fact that PH is at least as expressive as Weisfeiler-Lehman (WL) hierarchy [18], there recently emerged a series of works seeking to merge PH into graph neural networks (GNNs), delivering competitive performance on specific tasks [49, 18]. Standard schemes of existing works achieve this by employing pre-calculated topological features [49] or placing learnable filtration functions in the neural architectures [16, 18]. Such integration of PH features is claimed to enable GNNs to emphasize persistent topological sub-structures. However, it is still unclear to what extent the feature-level integration of PH is appropriate and how to empower GNNs with PH other than utilizing features.
@@
Graph pooling (GP) in parallel plays an important role in a series of graph learning methods [14], which hierarchically aggregates an upper-level graph into a more compact lower-level graph. Typically, GP relies on calculating an assignment matrix taking into account local structural properties such as community [32] and cuts [2]. Though the pooling paradigm in convolutional neural networks (CNNs) is quite successful [26], some researchers raise concerns about its effectiveness and applicability in graphs. For example, [29] challenges the local-preserving usage of GP by demonstrating that random pooling even leads to similar performance. Till now, it remains opaque what property should be preserved for pooled topology to better facilitate the downstream tasks.
@@
From Fig. 1(a), it is readily observed that PH and GP both seek to coarsen/sparsify a given graph in a hierarchical fashion: while PH gradually derives persistent sub-topology (substructures such as cycles) by adjusting the filtering parameter, GP obtains a sub-graph by performing a more aggressive cut-off. In a sense of understanding a graph through a hierarchical lens, PH and GP turn out to align with each other well.
@@
Driven by this observation, in this paper, we investigate the mechanism of aligning PH and GP so as to mutually reinforce each other. To this end, we conduct experiments by running a pioneer GP method DiffPool [47] to conduct graph classification on several datasets and at the same time use the technique in [16] to compute PH information. We manually change the pooling ratio and see what proportion of meaningful topological information (characterized by the ratio of non-zero persistence) is naturally preserved at the final training stage. Surprisingly, the correspondence is quite stable regardless of different datasets (see Fig. 1(b)), which implies the monotone trend between the pooling ratio and non-zero persistence is commonly shared by a large range of graph data. As a consequence, we develop a natural way to integrate PH and GP in both feature and topology levels. Concretely, in addition to concatenating vectorized PH diagram as supplementary features, we further enforce the coarsened graph to preserve topological information as much as possible with a specially designed PH-inspired loss function. Hence we term our method Topology-Invariant Pooling (TIP). TIP can be flexibly injected into a variety of existing GP methods, and demonstrates a consistent ability to provide substantial improvement over them. We summarize our contributions as follows:
@@ 2 Related work
Graph pooling. Graph pooling has been used in various applications, which can reduce the graph size while preserving its structural information. Early methods are based on clustering to coarsen graphs, such as the greedy clustering method Graclus [7], non-negative matrix factorization of the adjacency matrix [1], and spectral clustering [28]. Recently, learnable graph pooling methods have gained popularity, which learn to select important nodes in an end-to-end manner. DiffPool [47] follows a hierarchical learning structure by utilizing GNNs to learn clusters and gradually aggregate nodes into a coarser graph. MinCutPool [2] optimizes a normalized cut objective to partition graphs into clusters. DMoNPool [32] optimizes the modularity of graphs to ensure high-quality clusters. SEP [43] generates clusters in different hierarchies simultaneously without compromising local structures. These methods are classified as dense pooling due to the space complexity they incur. Despite their effectiveness, dense pooling methods have been criticized for high memory cost and complexity [5]. Therefore, various sparse pooling methods have been proposed, such as Top-K [11], ASAPool [36], and SAGPool [27]. These methods coarsen graphs by selecting a subset of nodes based on a ranking score. As they drop some nodes in the pooling process, these methods are criticized for their limited capacity to retain essential information, with potential effects on the expressiveness of preceding GNN layers [3].
Persistent homology in GNNs. PH is a technique to calculate topological features of structured data, and many approaches have been proposed to use PH in graph machine learning due to the high expressiveness of topological features on graphs [17]. Since non-isomorphic graphs may exhibit different topological features, the combination of PH and the Weisfeiler-Lehman (WL) algorithm leads to stronger expressive power [38]. This encourages further exploration on equipping GNNs with topological features. [49] propose that message passing in GNNs can be effectively reweighted using topological features. [16] and [18] provide theoretical and practical insights that filtraions in PH can be purely learnable, enabling flexible usage of topological features in GNNs. However, existing methods tend to view PH merely as a tool for providing supplementary information to GNNs, resulting in unsatisfactory improvements and limited interpretability.
@@ 3 Background
Let $\mathcal{G}=(V,E)$ be an undirected graph with $n$ nodes and $m$ edges, where $V$ and $E$ are the node and the edge sets, respectively. Nodes in attributed graphs are associated with features, and we denote by $V=\{(v,\mathbf{x}_{v})\}_{v\in 1:n}$ the set of nodes $v$ with $d$ dimensional attribute $\mathbf{x}_{v}$. It is also practical to represent the graph with an adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ and the node feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$.
@@
We focus on the general message-passing GNN framework that updates node representations by iteratively aggregating information from neighbors [12]. Concretely, the $k$-th layer of such GNNs can be expressed as:
@@
where $\mathbf{\theta}^{(k)}$ is the trainable parameter, and $\mathrm{M}$ is the message propagation function. Numbers of $\mathrm{M}$ have been proposed in previous research [24, 15]. A complete GNN is typically instantiated by stacking multiple layers of Eq. 1. Hereafter we denote by $\mathrm{GNN}(\cdot)$ an arbitrary such multi-layer GNN for brevity.
@@
GP in GNNs is a special layer designated to produce a coarsened or sparsified sub-graph. Formally, GP can be formulated as $\mathcal{G}\mapsto\mathcal{G}_{P}=(V_{P},E_{P})$ such that the number of nodes $|V_{P}|\leq n$. GP layers can be placed into GNNs in a hierarchical fashion to persistently coarsen the graph. Typical GP approaches [47, 2, 32] rely on learning a soft cluster assignment matrix $\mathbf{S}^{(l)}\in\mathbb{R}^{n_{l-1}\times n_{l}}$:
@@
These approaches differ from each other in the way to produce $\mathbf{S}$, which is used to inject a bias in the formation of clusters. In our work, we select three GP methods, i.e., DiffPool [47], MinCutPool [2], and DMoNPool [32], to cope with. Details of these methods are in Appendix A.
@@
Simplicial complexes stand as the focal point within the realm of algebraic topology. An assembly of simplices of certain dimensions constitutes a simplicial complex denoted as $K$. A graph can be seen as a low-dimensional simplicial complex that only contains 0-simplices (vertices) and 1-simplices (edges) [18]. The simplest kind of topological features describing graphs are Betti numbers, formally denoted as $\beta_{0}$ for the number of connected components and $\beta_{1}$ for the number of cycles.
Despite the limited expressive power of these two numbers, it can be improved by evaluating them alongside a filtration. Filtrations are scalar-valued functions of the form $f:V\cup E\rightarrow\mathbb{R}$ , which assigns each vertex and edge a value. Changes in the Betti numbers, named as persistent Betti numbers, can subsequently be monitored throughout the progress of the filtration: by considering a threshold $(a\in\mathbb{R})$, we can analyze the subgraph originating from the pre-image of $((-\infty,a])$ of $f$, denoted as $(f^{-1}((-\infty,a]))$. The image of $f$ leads to a finite set of values $a_{1}<\cdots<a_{n}$ and generates a sequence of nested subgraphs of the form $\emptyset\subseteq\mathcal{G}_{0}\subseteq\ldots\mathcal{G}_{k}\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$, where $\mathcal{G}_{k}=(V_{k},E_{k})$ is a subgraph of $\mathcal{G}$ with $V_{k}:=\left\{v\in V\mid f\left(\mathbf{x}_{v}\right)\leq a_{k}\right\}$ and $E_{k}:=\left\{(v,w)\in E\mid\max\left\{f\left(x_{v}\right),f\left(x_{w}\right)\right\}\leq a_{k}\right\}$. This process is also known as persistent homology (denoted as $ph(\cdot)$) on graphs. Typically, persistent Betti numbers are summarized in a persistence diagram (PD) as $\mathrm{ph}(\mathcal{G},f)=\{\mathcal{D}_{0},\mathcal{D}_{1},...\}$, which is made up of tuples $(a_{i},a_{j})\in\mathbb{R}^{2}$, with $a_{i}$ and $a_{j}$ representing the creation and destruction of a topological feature respectively (see Fig. 1(c)). The absolute difference in function values $|a_{j}-a_{i}|$ is called the persistence of a topological feature, where high persistence corresponds to features of the function, while low persistence is typically considered as noise [18, 37].
@@ 4.1 Overview
Figure 2: Overview of our method. The shaded part is one layer of Topology-Invariant Pooling.
@@ 4.2 Topology-Invariant Pooling
In many real-world applications, the topology of graphs are of utmost importance [41, 46, 16]. However, typical GNNs fail to capture certain topological structures in graphs, such as cycles [4, 48, 20]. Moreover, in dense graph pooling, graphs are pooled without preserving any topology. Even if we manage to make GNN topology-aware, the coarsened graph is nearly fully connected and has no meaningful topology at all, impairing the use of GNNs in these tasks. To overcome these limitations, we propose to inject topological information into GP. We resort to PH to characterize the importance of edges.
The core of PH is the notion of filtration, which presents a challenging task to choose the right filtration. As the coarsened graph evolves in each training step, integrating PH into GP demands multiple computations of filtrations. To address this, we place recently proposed learnable filtration (LF) functions [16] to incorporate PH information for flexibility and efficiency. LF relies on node features and graph topology, which are readily available in GP. Consequently, LF can be seamlessly integrated into GP with minimal computational overhead. Specifically, we employ an MLP network $\Phi(\cdot)$ as the filtration function together with $\mathrm{sigmoid}(\cdot)$ to map node features $\mathbf{X}\in\mathbb{R}^{n\times d}$ into $n$ scalar values. Recently, an increasing amount of attention has been devoted to cycles [4, 48, 20] due to their significant relevance to downstream tasks in various domains such as biology [25], chemistry [33], and social network analysis [22]. Recognizing that cycles offer an intuitive representation of graph structure, we focus on the one-dimensional PD associated with cycles. For those edges do not form cycles, their creation and destruction are the same, leading to zero persistence. Following the standard way in GP (Eq. 2 3 4), we additionally propose the subsequent modules to inject PH into GP at both feature and topology levels.
@@
One major limitation of utilizing LF proposed in [16] is that the computation process is unaware of edge weights, i.e. edges with non-zero weights will be treated equally, so PH cannot directly extract meaningful topology from $\mathbf{A}^{(l)}$. Besides, rethinking GP in Eq. 3, the coarsened adjacency matrix has limited expressive power for two reasons. First, although $\mathbf{S}^{(l)}$ is a soft assignment matrix obtained by $\mathrm{softmax}(\cdot)$, each element still has nonzero values, i.e. $\mathbf{A}^{(l)}$ is always densely connected. Second, the edge weights may span a wide range by multiplication (see Appendix D for empirical evidence). These drawbacks hinder the stability and generalization power of the subsequent message passing layers [13]. None of the existing GP methods can handle these problems properly.
@@
where $\mathbf{A}^{(l)}$ is first normalized in the range of $[0,1]$, and $\mathrm{resample}(\cdot)$ is performed independently for each matrix entry using the Gumbel-softmax trick [21]. In practice, only the upper triangular matrix is resampled to make it symmetric and we add self-loops to the graph.
@@
The aforementioned mechanism can explicitly inject topological information into graphs, but it relies on the condition that the coarsened graph retains certain essential sub-topology. To this end, we propose an additional loss function to guide the GP process. Intuitively, the coarsened graph should exhibit similarity to the original graph in terms of topology. Since the computation of PH is differentiable, one possible approach is to directly minimize the differences between the PDs of the original graph and the coarsened graph. However, this implementation would require computing the Wasserstein distance between two PDs through optimal transport [45], which is intractable in training due to its complexity. Considering that our objective is to estimate the difference, we instead propose vectorizing the PDs and minimizing their high-order statistical features [34]. Specifically, we use several transformations (denoted as $\mathrm{transform(\cdot)}$) and concatenate the output, including triangle point transformation, Gaussian point transformation and line point transformation introduced in [6] to convert the tuples in PD into vector $\mathbf{h}_{t}$ $(t\in[1,m])$. We calculate the mean vector $\mathbf{\mu}$ as well as the second-order statistics as the standard deviation vector $\mathbf{\sigma}$ as:
$\mathbf{h}_{t}=\mathrm{transform(\mathcal{D}_{1})}$ (7)